What volume of a solution of potassium permanganate ( 81 mg / mL) is required for the preparation of 8.3 L of a solution containing the substance at a concentration of 1 in 4,000?(no decimal place)
Answer #1

Potassium permanganate 81 mg in 1ml =1000L

How much concentration of potassium permanganate is required for 8.3 L

Using formula : C1 V1 = C2 V2

C1 is 81mg

V1 = 1000L

C2 = ?

V2 = 8.3L

=> 81×1000= C2 ×8.3

C2 = 81×1000 / 8.3

C2 = 9759mg